# Institutionalizing Folk Theories of Algorithms: How Multi-Channel Networks (MCNs) Govern Algorithmic Labor in Chinese Live-Streaming Industry

### v1

@@ 6.2. Folk Theories of Algorithms for Labor Management
While most scholarship on folk theories emphasized their adaptive or empowering qualities, helping users make sense of opaque systems in the absence of formal knowledge (Huang et al., 2022; Siles et al., 2020; Bonini and Treré, 2024; Eslami et al., 2016), our findings suggest a more complex and strategic function: folk theories are not only interpretive frameworks, but also governance tools. In the context of live-streaming platforms, they become a key mechanism through which institutions like MCNs direct labor, manage uncertainty, and externalize risk. This repositions folk algorithmic theories from being user-centric artifacts to organizational instruments of soft control.
@@
A growing body of CSCW and HCI literature has investigated the dynamics of algorithmic management, especially in gig economies and platform labor systems (Kellogg et al., 2020; Möhlmann et al., 2021). Here, scholars have shown how algorithms allocate work, evaluate performance, and discipline workers without the need for traditional human supervision. However, this line of work often focuses on direct algorithmic feedback (e.g., ratings, scores, rankings) or automated decision-making (e.g., job assignments). Our findings build on this by revealing a second layer of algorithmic management: the interpretive layer, where algorithmic outputs are not simply accepted or resisted, but actively narrated, framed, and naturalized through folk theories constructed by intermediaries. These folk theories are not mere byproducts of algorithmic opacity but are engineered and circulated by MCNs to align with institutional interests and govern labor at scale.
Given the opacity of platform algorithms  (Ma et al., 2025; Deng et al., 2023b; Nagaraj Rao et al., 2025), most workers lack access to how algorithmic decisions are made. While often seen as a limitation, in our case, opacity can also be leveraged as strategic ambiguity  (Gillespie, 2014; Ananny and Crawford, 2018; Eubanks, 2018). Our study shows how MCNs capitalize on this: in the absence of verified knowledge, they develop external simplified theories of algorithmic causality for streamers to set expectations, justify pressure, and shape behavior, without guaranteeing outcomes.
These external folk theories function as proxies for algorithmic authority. They tell streamers what the algorithm “wants” and how to “win” it over, but the logic they offer is shaped by MCNs’ managerial goals. By promoting these ideas as common sense, MCNs shift responsibility for failure onto streamers. Poor outcomes are framed as personal shortcomings, such as insufficient effort, investment, or affect, rather than algorithmic opacity.
Notably, this mode of governance does not depend on technical accuracy. As our findings show, even MCN staff admit their advice is speculative. But as long as these external theories appear confident, actionable and correlate with visible metrics, they retain influence to streamers. This performativity echoes insights from critical algorithm studies  (Gillespie, 2014; Eubanks, 2018): sometimes belief in algorithmic logic shapes behavior, regardless of its truth.
@@
MCNs therefore construct what we call interpretive infrastructures—a system of informal yet durable knowledge that governs through belief, not verification/reality. Our findings contribute to CSCW’s growing attention to the epistemic and organizational dimensions of algorithmic labor. Folk theories, we argue, are not peripheral but are central to how algorithmic management operates in opaque, high-risk environments. They enable institutions to govern not through direct technical means, but through the circulation of strategically engineered narratives. This invites future research into how informal knowledge can become a durable, institutionally mediated mode of control in algorithmic workplaces.
@@ 6.3. Design and Policy Implications
Finally, our work invites CSCW researchers, designers, and policymakers to view algorithmic sensemaking not as an individual cognitive task but as an organizational process. Institutions do not merely allocate labor or payment, and mediate access to meaning. Rather than only designing tools to increase algorithmic transparency, we invite CSCW and HCI scholars to ask: who gets to construct algorithmic meaning, for whom, and to what ends? Designing technologies for algorithmic collaboration therefore requires grappling with the political economy of epistemic authority and these questions raise broader design and policy challenges for shaping the future of algorithmic work.
@@ 7. Limitations and Future Work
As with all qualitative research, our findings are shaped by the specific contexts, participants, and interpretive lenses of our study. There are a few important limitations. First, this research draws on ethnographic fieldwork conducted in three MCNs in China, and while we prioritized diversity in location, streamer profile, and organizational structure, our sample cannot represent the full heterogeneity of the live-streaming industry, especially given regional, platform-specific, and organizational variation. Second, our focus on China brings unique sociopolitical and economic conditions that influence how platform labor and algorithmic governance unfold, including platform regulation, labor precarity, and the prominence of MCNs. While these conditions offer valuable insights into the global dynamics of platform intermediation, they may not be directly generalizable to other contexts. Third, our access to internal MCN practices was inevitably limited: despite prolonged immersion and trust-building, some dimensions of organizational decision-making and platform negotiation remained opaque or inaccessible. Finally, our analysis focuses on the interpretive and discursive construction of algorithmic knowledge, but does not directly evaluate the technical validity of the folk theories themselves or track their longitudinal impact on streamer outcomes – areas that future mixed-method or longitudinal studies could explore.
@@ 8. Conclusion
This study examined how algorithmic labor is governed not solely through platform design, but through the interpretive and organizational practices of intermediary institutions. Rather than focusing only on platform power or individual resistance, our analysis highlights how Multi-Channel Networks (MCNs) in the Chinese live-streaming industry play a central role in constructing and institutionalizing folk theories of algorithms. We identify a system of dual algorithmic theories: internally, MCNs acknowledge the volatility and opacity of platform systems, adopting probabilistic strategies to manage organizational risk; externally, they promote simplified, prescriptive narratives that portray the algorithm as transparent, fair, and responsive to individual effort. We show that these folk theories, though speculative, unverifiable and often inaccurate, can be mobilized as tools of labor control: providing streamers with simplified theories of algorithmic success, converting platform uncertainty into actionable scripts, and shifting responsibility for failure onto individuals. We call for future research to take seriously the role of intermediary institutions in algorithmic management and to attend to how folk theories of algorithms, when institutionalized, can shape the possibilities and limits of platform labor. By tracing how folk theories are used not only to explain uncertain algorithms, but to manage platform labor, this study reveals how interpretation itself becomes a key site of power in the algorithmic age.

### v2

@@ 7.2. Discursive Construction of Algorithms for Labor Management
While most scholarship on algorithmic interpretations emphasized their adaptive or empowering qualities, helping users make sense of opaque systems in the absence of formal knowledge (Huang et al., 2022; Siles et al., 2020; Bonini and Treré, 2024; Eslami et al., 2016), our findings suggest a more complex and strategic function: algorithmic interpretations are not only interpretive frameworks, but also labor governance tools. In the context of live-streaming platforms, they become a key mechanism through which institutions like MCNs direct labor, manage uncertainty, and externalize risk.
@@
A growing body of CSCW and HCI literature has investigated the dynamics of algorithmic management, especially in gig economies and platform labor systems (Kellogg et al., 2020; Möhlmann et al., 2021). Here, scholars have shown how algorithms allocate work, evaluate performance, and discipline workers without the need for traditional human supervision. However, this line of work often focuses on direct algorithmic feedback (e.g., ratings, scores, rankings) or automated decision-making (e.g., job assignments). Our findings build on this by revealing a second layer of algorithmic management: the interpretive layer, where algorithmic outputs are not simply accepted or resisted, but actively narrated, framed, and naturalized through algorithmic interpretations constructed by intermediaries. These algorithmic interpretations are not mere byproducts of algorithmic opacity but are engineered and circulated by MCNs to align with institutional interests and govern labor at scale.
Central to this mode of labor governance is a pronounced epistemic asymmetry between MCNs and live-streamers. MCNs occupy a privileged position in the circulation of algorithmic knowledge: they control when, how, and in what form explanations about algorithmic behavior are made available to streamers. Although MCNs do not possess verified or complete knowledge of platform algorithms, they nevertheless monopolize interpretive authority by presenting themselves as the most credible source of explanation in an otherwise opaque system. Streamers, by contrast, typically lack the resources, comparative data, and temporal distance needed to independently evaluate algorithmic claims or to develop alternative explanatory frameworks.
This asymmetry is not merely informational but organizationally produced. MCNs’ internal recognition of algorithmic uncertainty remains largely inaccessible to streamers, while externally circulated interpretations are selectively simplified, stabilized, and framed as actionable truths. As a result, streamers encounter algorithmic outcomes through a narrowed interpretive channel, where uncertainty is individualized and responsibility is localized. Even when outcomes fail to materialize, streamers often lack the epistemic grounds to contest MCNs’ explanations, leading doubt to remain fragmented and personalized rather than collective or structural.
Through this asymmetry, MCNs are able to convert limited and speculative knowledge into durable authority. Control is exercised not by withholding information entirely, but by curating the forms of knowledge that circulate and defining what counts as legitimate understanding of algorithmic success and failure. In this sense, epistemic asymmetry becomes a key condition for the effectiveness of discursive construction: it allows algorithmic interpretations to function as governance tools precisely because alternative interpretations are difficult to verify, share, or stabilize among workers.
@@
MCNs therefore construct what we call interpretive infrastructures—a system of informal yet durable knowledge that governs through belief, not verification/reality. Our findings contribute to CSCW’s growing attention to the epistemic and organizational dimensions of algorithmic labor. Algorithmic interpretations, we argue, are not peripheral but are central to how algorithmic management operates in opaque, high-risk environments. They enable institutions to govern through the circulation of strategically engineered narratives. This invites future research into how informal knowledge can become a durable, institutionally mediated mode of control in algorithmic workplaces.
Additionally, a deeper exploration of the dynamic between external and internal algorithmic interpretations would also enrich this discussion around labor management. In our study, external algorithmic interpretations constructed by MCNs serve managerial interests, stabilizing labor through prescriptive narratives. In contrast, recent work highlights how individual users develop their own folk theories as diagnostic tools for navigating platform logics and adapting their own practices (Mayworm et al., 2024). This comparison underscores two divergent trajectories: organizational actors such as MCNs strategically engineer algorithmic interpretations to manage others’ labor, while users mobilize folk theories to make opaque systems actionable for themselves. Future inquiries must therefore pay closer attention to whose interests these algorithmic interpretations represent and what purposes they serve. Importantly, algorithmic interpretations should not be assumed as inherently empowering or exploitative; rather, they occupy an ambivalent position that can be instrumentalized in different directions. For labor researchers, unions, and advocates, this means recognizing that algorithmic interpretations are contingent tools whose effects depend on the power relations and institutional agendas through which they are mobilized.
@@ 7.3. Design and Policy Implications
Finally, our work invites CSCW researchers, designers, and policymakers to view algorithmic interpretations as an organizational labor management process. Institutions do not merely allocate labor or payment, and mediate access to meaning. Rather than only designing tools to increase algorithmic transparency, we invite CSCW and HCI scholars to ask: who gets to construct algorithmic meaning, for whom, and to what ends? Designing technologies for algorithmic collaboration therefore requires grappling with the political economy of epistemic authority and these questions raise broader design and policy challenges for shaping the future of algorithmic work.
@@ 8. Limitations and Future Work
As with all qualitative research, our findings are shaped by the specific contexts, participants, and interpretive lenses of our study. There are a few important limitations. First, this research draws on ethnographic fieldwork conducted in three MCNs in China, and while we prioritized diversity in location, streamer profile, and organizational structure, our sample cannot represent the full heterogeneity of the live-streaming industry. Second, our focus on China brings unique sociopolitical and economic conditions that influence how platform labor and algorithmic governance unfold. While these conditions offer valuable insights into the global dynamics of platform intermediation, they may not be directly generalizable to other contexts. Third, our access to internal MCN practices was inevitably limited: despite prolonged immersion and trust-building, some dimensions of organizational decision-making and platform negotiation remained opaque. Relatedly, our study foregrounds epistemic asymmetry but does not systematically examine how such asymmetries vary across organizational settings or how they might be mitigated in practice. Future research could investigate the conditions under which interpretive authority over algorithms is concentrated or redistributed, as well as the implications of these dynamics for workers’ agency. Finally, our analysis focuses on the discursive construction of algorithms, but does not directly evaluate the technical validity of algorithmic interpretations themselves or track their longitudinal impact on streamer outcomes—areas that future mixed-method studies could explore.
@@ 9. Conclusion
This study examined how algorithmic labor is governed through the interpretive and organizational practices of intermediary institutions besides platforms themselves. Rather than focusing only on platform power or individual behaviors, our analysis highlights how Multi-Channel Networks (MCNs) in Chinese live-streaming industry play a central role in constructing algorithmic interpretations. We identify a system of dual algorithmic interpretations: internally, MCNs acknowledge the volatility and opacity of platform systems, adopting probabilistic strategies to manage organizational risk; externally, they promote simplified, prescriptive narratives that portray the algorithm as transparent, fair, and responsive to individual effort. We argue that these interpretations, though speculative, unverifiable, and often inaccurate, can be mobilized as tools of labor control: providing streamers with simplified interpretations of algorithmic success, converting platform uncertainty into actionable scripts, and shifting responsibility for failure onto individuals. We call for future research to focus on the role of intermediary institutions in algorithmic management and to attend to how discursive construction of algorithms shapes the possibilities and limits of platform labor. By tracing how algorithmic interpretations are used not only to explain uncertain algorithms, but to manage platform labor, this study reveals how interpretation itself becomes a key site of power in the algorithmic age.